OpenStudy (anonymous):

how to solve h^2-5h-36=0

OpenStudy (anonymous):

first of all find discriminant

OpenStudy (anonymous):

-36 = -9*4 -5 = -9+4 So, the factors are (h-9)(h+4)

OpenStudy (anonymous):

ohh wow..i be doing it wrong..

OpenStudy (anonymous):

No medal??

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

this question is too easy to get medal isn't it? :D

OpenStudy (anonymous):

hey man....whatever gets it! :-)